# Sparse Hanson-Wright Inequalities with Applications

### v3

@@ 6.1 Proof of Theorem 4
    and $u=\log[(r/\Delta)^{\frac{1}{\beta}}]>\frac{1}{\beta}>0$. The inequality becomes $e^{u}\leq Lu.$ By Proposition 1, the feasibility of this inequality implies that $L=\beta T^{\frac{1}{\beta}}>e$ and its solution is
    $-u\geq W_{-1}(-L^{-1}),$
@@ 6.2 Proof of Theorem 5
    Note that by the supposition $r/\Delta_{1}>e$ and $\beta\in(0,1]$, $u>1$. By Proposition 1, when $L_{1}>e$, (6.25) is solvable and its solution is
@@
    which, by Proposition 1, is solvable when $L_{2}>e$ and we have
@@
We optimize $g_{\alpha}(k,t)$ for $21\leq t\leq k/$ 2 and $1\leq k\leq r$ by considering two cases.
@@
Since $r^{3/2}/\lambda_{2}>1$, using the same argument as in Case 2 of the proof of Lemma 6.1, we optimize $\log g_{\alpha}(k,k/2)$ and obtain
$g_{\alpha}(k,t)\leq k_{*}^{(2-\frac{2}{\alpha})r}\quad\text{with}\quad k_{*}=\left(2-\frac{2}{\alpha}\right)\frac{r}{\log(r^{3/2}/\Delta_{1})}$
@@ 7.1 Proof of Theorem 6
To derive (7.2), we first observe that $\|X_{i}^{2}-\mathbb{E}X_{i}^{2}\|_{L_{r}}\leq 2\|X_{i}^{2}\|_{L_{r}}\leq(p_{i}2^{r}K^{2r}r^{\frac{2}{\alpha}})^{1/r}$. Then we apply Theorem 2 to the centered random variables $X_{i}^{2}-\mathbb{E}X_{i}^{2}$ to obtain (7.2). Without loss of generality, we assume $C_{\alpha}\geq 1$.
@@
Let us assume Theorem 11 holds and prove Theorem 6. Given $D>0$ and $\delta\in(0,1)$, we take $r=\log n$, $t=\exp(\frac{D+7}{2})$. Let $\mathtt{f}(D,\delta):=\max\{1,\frac{(20\delta)^{4}}{12^{4}}\}.$ Then the conditions of Theorem 11 are satisfied provided that
$p\geq\frac{12^{6}e^{2(D+7)}\mathtt{f}(D,\delta)}{\delta^{4}}\cdot\frac{C_{\alpha}^{2}K^{4}(\log n)^{\max\{\frac{2}{\alpha},2\}}}{n}$
@@
$\eta\geq\frac{12^{6}e^{2(D+7)}\mathtt{f}(D,\delta)}{\delta^{4}}\cdot\frac{C_{\alpha}^{2}K^{4}(\log n)^{2}}{n}.$
@@
$L(\delta,D):=\frac{12^{6}e^{2(D+7)}\mathtt{f}(D,\delta)}{\delta^{4}},$ (7.5)
@@
where the matrix $G_{0}^{(i)}$ is obtained from $G^{(i)}$ by setting all diagonal entries to zero. Observe that $\phi^{(i)}\|G_{0}^{(i)}\|_{\max}\leq 2$ and by Cauchy-Schwartz inequality and (7.9), we have
@@ Acknowledgments
We thank Fanny Augeri, Antti Knowles, Stanislav Minsker, Roman Vershynin, and Shuheng Zhou for helpful discussions. K. Wang is partially supported by Hong Kong RGC grant GRF 16304222 and ECS 26304920. Y. Zhu was partially supported by NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning and an AMS-Simons Travel Grant. This material is based upon work supported by the Swedish Research Council under grant no. 2021-06594 while Y. Zhu was in residence at Institut Mittag-Leffler in Djursholm, Sweden during the Fall of 2024.
@@ A.1 Auxiliary lemmas
Note that the function $f(x)=(x+1)(1+\frac{1}{x})^{x}$ is strictly in creasing on $(0,+\infty)$. We know the inequality holds as $f(a-1)<f(b)$.
@@ A.2 The Lambert W function
has real solutions for $y$ only when $x\geq-e^{-1}$. When $x\geq 0$, there exists a unique solution given by the principal branch of the Lambert W function, denoted as $y=W_{0}(x)$. For $-e^{-1}\leq x<0$, there are exactly two solutions: the principal branch $y=W_{0}(x)$ and the lower branch $y=W_{-1}(x)$. The two branches of the Lambert W function are illustrated in Figure 1. We include a result about the Lambert W function that is used in our proof.
@@
This inequality (A.2) is solvable only when $-L^{-1}\geq-e^{-1}$ since the minimum of $xe^{x}$ is $-e^{-1}$. Specially, if $L=e$, then $u=1$. For $L\geq e$, $-L^{-1}\in[-e^{-1},0)$ and from the definition and behavior of Lambert W function (see Figure 1), we find that $W_{-1}(-L^{-1})\leq-u\leq W_{0}(-L^{-1})$. This completes the proof. ∎
@@ B.1 Proof of Theorem 8
Since $\check{H}$ has independent bounded entries, [44, Theorem 4.8] implies
@@
which yields $|u_{i}(k)|\leq\sqrt{\frac{2}{\eta}}=\sqrt{2L}C_{\alpha}K^{2}\frac{\log n}{\sqrt{n}}$. This completes the proof.
@@ B.2 Proof of Theorem 9
For $\alpha\geq 2$, we provide a stronger spectral norm concentration bound. The proof is based on a general result from [18] for the spectral norm of inhomogeneous random matrices where each entry has a symmetric distribution. We apply a symmetrization argument to obtain the following bound:
@@
Let $X^{\prime}$ be a symmetric random matrix such that $X^{\prime}_{ij}=\varepsilon_{ij}X_{ij}$, where $\varepsilon_{ij}$ are i.i.d. Rademacher random variables. Then $X^{\prime}_{ij}$ has a symmetric distribution. From [59, Exercise 6.4.5] for any $t\geq 0$, $\mathbb{E}e^{t\|X\|}\leq\mathbb{E}e^{2t\|X^{\prime}\|}$. From [18, Lemma 3.17], assuming $\sigma_{*}=1$, and let $d=\lceil\sigma^{2}\rceil,$ we have $\mathbb{E}\operatorname{tr}e^{tX^{\prime}}\leq e^{2\sqrt{d}t+t^{2}}$. Since $e^{t\|X^{\prime}\|}\leq n\operatorname{tr}e^{tX^{\prime}}+n\operatorname{tr}e^{-tX^{\prime}}$, we obtain $\mathbb{E}[e^{t\|X^{\prime}\|}]\leq 2ne^{2\sqrt{d}t+t^{2}}$. Then $\mathbb{E}e^{t\|X\|}\leq 2ne^{4\sqrt{d}t+4t^{2}}$. By Markov’s inequality, for any $\varepsilon\geq 0$,
@@
Conditioned on the event :=$\mathcal{E}:=\{\max_{i}\sum_{j}X_{ij}\leq C_{2}np\}$, by choosing $C_{2}$ sufficiently large, we can apply Lemma B.2 to $\frac{1}{CK}\tilde{H}$ to get
@@
The proof of Theorem 9 follows a similar approach to that of Theorem 8. The key difference lies in the application of improved moment estimates in Lemma 6.2 and Lemma 6.4. For brevity, we present only the essential estimates, omitting the full derivation.
@@
For $k\geq 1$, we show that conditioning on $\Omega_{k-1}\cap\Xi_{k-1}$, the event $\Omega_{k}\cap\Xi_{k}$ holds with high probability. Define a threshold index $\widetilde{\mathcal{K}}:=\min\{k\in\mathbb{N}:k\leq\mathcal{K},|m(z_{k})-\tilde{m}(z_{k})|\leq 5\zeta_{k}t\}.$ For $k<\widetilde{\mathcal{K}}$, $|m(z_{k})-\tilde{m}(z_{k})|>5\zeta_{k}t$, and we must sure that $s(z_{k})$ is close to $m(z_{k})$ rather than $\tilde{m}(z_{k})$. In contrast, for $k\geq\widetilde{\mathcal{K}}$, the distinction between $m(z_{k})$ and $\tilde{m}(z_{k})$ becomes insignificant, leading to a simpler argument. The majority of the proof follows steps analogous to those used for the base case $k=0$. We provide a sketch of the key arguments below.
@@ D.1 Proof of Corollary 1
$\mathbb{E}|X_{i}^{2}-\mathbb{E}X_{i}^{2}|^{l}\leq 2^{l}\mathbb{E}|X_{i}|^{2l}\leq 2^{l}p_{i}K^{2l}(2l)^{\frac{2l}{\alpha}}.$ (D.1)

### v4

@@ 6.1 Proof of Theorem 9
    and $u=\log[(r/\Delta)^{\frac{1}{\beta}}]>\frac{1}{\beta}>0$. The inequality becomes $e^{u}\leq Lu.$ By Proposition A.1, the feasibility of this inequality implies that $L=\beta T^{\frac{1}{\beta}}>e$ and its solution is
    $u\geq W_{-1}(-L^{-1}),$
@@ 6.2 Proof of Theorem 10
    Note that by the supposition $r/\Delta_{1}>e$ and $\beta\in(0,1]$, $u>1$. By Proposition A.1, when $L_{1}>e$, (6.30) is solvable and its solution is
@@
    which, by Proposition A.1, is solvable when $L_{2}>e$ and we have
@@
We optimize $g_{\alpha}(k,t)$ for $1\leq t\leq k/2$ and $1\leq k\leq r$ by considering two cases.
@@
Since $r^{3/2}/\Delta_{2}>1$, using the same argument as in Case 2 of the proof of Lemma 6.1, we optimize $\log g_{\alpha}(k,k/2)$ and obtain
$g_{\alpha}(k,t)\leq k_{*}^{(2-\frac{2}{\alpha})r}\quad\text{with}\quad k_{*}=\left(2-\frac{2}{\alpha}\right)\frac{r}{\log(r^{3/2}/\Delta_{2})}$
@@ 7.1 Proof of Theorem 4
To derive (7.2), we first observe that $\|X_{i}^{2}-\mathbb{E}X_{i}^{2}\|_{L_{r}}\leq 2\|X_{i}^{2}\|_{L_{r}}\leq(p_{i}2^{r}K^{2r}r^{\frac{2}{\alpha}})^{1/r}$. Then we apply Theorem 2 to the centered random variables $X_{i}^{2}-\mathbb{E}X_{i}^{2}$ to obtain (7.2). Without loss of generality, we assume $C_{\alpha}\geq 1$. Recall that $\widetilde{H}$ is the normalized version of $H$ whose entries satisfy Assumption 1.
@@
Let us assume Theorem 11 holds and prove Theorem 4. Given $D>0$ and $\delta\in(0,1)$, we take $r=\log n$, $t=\exp(\frac{D+7}{2})$. Let $\mathtt{f}(\delta):=\max\{1,\frac{(20\delta)^{4}}{12^{4}}\}.$ Then the conditions of Theorem 11 are satisfied provided that
$p\geq\frac{12^{6}e^{2(D+7)}\mathtt{f}(\delta)}{\delta^{4}}\cdot\frac{C_{\alpha}^{2}K^{4}(\log n)^{\max\{\frac{2}{\alpha},2\}}}{n}$
@@
$\eta\geq\frac{12^{6}e^{2(D+7)}\mathtt{f}(\delta)}{\delta^{4}}\cdot\frac{C_{\alpha}^{2}K^{4}(\log n)^{2}}{n}.$
@@
$L(\delta,D):=\frac{12^{6}e^{2(D+7)}\mathtt{f}(\delta)}{\delta^{4}},$ (7.5)
@@
where the matrix $G_{0}^{(i)}$ is obtained from $G^{(i)}$ by setting all diagonal entries to zero. Observe that $\phi^{(i)}\|G_{0}^{(i)}\|_{\max}\leq 2$ and by Cauchy-Schwarz inequality and (7.9), we have
@@ Acknowledgments
We thank Fanny Augeri, Antti Knowles, Stanislav Minsker, Roman Vershynin, and Shuheng Zhou for helpful discussions. K. Wang is partially supported by Hong Kong RGC grant GRF 16304222 and ECS 26304920. Y. Zhu was partially supported by NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning, an AMS-Simons Travel Grant, and the Simons Grant MPS-TSM-00013944. This material is based upon work supported by the Swedish Research Council under grant no. 2021-06594 while Y. Zhu was in residence at Institut Mittag-Leffler in Djursholm, Sweden during the Fall of 2024.
@@ A.1 Auxiliary lemmas
Note that the function $f(x)=(x+1)(1+\frac{1}{x})^{x}$ is strictly in creasing on $(0,+\infty)$. We know the inequality holds as f(a1)<f(b)f(a-1)<f(b).
@@ A.2 The Lambert W function
has real solutions for $y$ only when $x\geq-e^{-1}$. When $x\geq 0$, there exists a unique solution given by the principal branch of the Lambert W function, denoted as $y=W_{0}(x)$. For $e^{-1}\leq x<0$, there are exactly two solutions: the principal branch $y=W_{0}(x)$ and the lower branch $y=W_{-1}(x)$. The two branches of the Lambert W function are illustrated in Figure 1. We include a result about the Lambert W function that is used in our proof.
@@
This inequality (A.2) is solvable only when $L^{-1}\geq-e^{-1}$ since the minimum of $xe^{x}$ is $e^{-1}$. Specially, if $L=e$, then $u=1$. For $L\geq e$, $L^{-1}\in[-e^{-1},0)$ and from the definition and behavior of Lambert W function (see Figure 1), we find that $W_{-1}(-L^{-1})\leq-u\leq W_{0}(-L^{-1})$. This completes the proof. ∎
@@ B.1 Proof of Theorem 6
Since $\check{H}$ has independent bounded entries, [47, Theorem 4.8] implies
@@
which yields $|u_{i}(k)|\leq\sqrt{{2}{\eta}}=\sqrt{2L}C_{\alpha}K^{2}\frac{\log n}{\sqrt{n}}$. This completes the proof.
@@ B.2 Proof of Theorem 7
For $\alpha\geq 2$, we provide a stronger spectral norm concentration bound. The proof is based on a general result from [19] for the spectral norm of inhomogeneous random matrices, where each entry has a symmetric distribution. We apply a symmetrization argument to obtain the following bound:
@@
Let $X^{\prime}$ be a symmetric random matrix such that $X^{\prime}_{ij}=\varepsilon_{ij}X_{ij}$, where $\varepsilon_{ij}$ are i.i.d. Rademacher random variables. Then $X^{\prime}_{ij}$ has a symmetric distribution. From [65, Exercise 6.4.5] for any $t\geq 0$, $\mathbb{E}e^{t\|X\|}\leq\mathbb{E}e^{2t\|X^{\prime}\|}$. From [19, Lemma 3.17], assuming $\sigma_{*}=1$, and let $d=\lceil\sigma^{2}\rceil,$ we have $\mathbb{E}\operatorname{tr}e^{tX^{\prime}}\leq e^{2\sqrt{d}t+t^{2}}$. Since $e^{t\|X^{\prime}\|}\leq n\operatorname{tr}e^{tX^{\prime}}+n\operatorname{tr}e^{-tX^{\prime}}$, we obtain $\mathbb{E}[e^{t\|X^{\prime}\|}]\leq 2ne^{2\sqrt{d}t+t^{2}}$. Then $\mathbb{E}e^{t\|X\|}\leq 2ne^{4\sqrt{d}t+4t^{2}}$. By Markov’s inequality, for any $\varepsilon\geq 0$,
@@
Conditioned on the event $\mathcal{E}:=\{\max_{i}\sum_{j}X_{ij}\leq C_{2}np\}$, by choosing $C_{2}$ sufficiently large, we can apply Lemma B.2 to $\frac{1}{CK}\tilde{H}$ to get
@@
The proof of Theorem 7 follows a similar approach to that of Theorem 6. The proof of Theorem 7 follows a similar approach to that of Theorem 6; the key difference is that, while Theorem 6 uses only the baseline concentration bounds (Theorems 1 and 2), the proof of Theorem 7 additionally relies on the Bennett-type refinements (Theorems 9 and 10) via the improved moment estimates in Lemmas 6.2 and 6.4, which are sharper in the large-deviation regime. Concretely, these improved inputs yield an additional logarithmic gain (visible through the $\log(\cdot)]^{-\beta}$ factors in Lemma B.4), which strengthens the large-deviation error parameters and reduces the sparsity threshold needed to close the self-consistent estimates from $p\gtrsim\log^{2}n/n$ (the regime of Theorem 6) to $p\gtrsim\log n/n$, enabling Theorem 7. For brevity, we present only the essential estimates, omitting the full derivation.
@@
By Vieta’s formula, $m(z)+\tilde{m}(z)=-z$. Therefore, for $z_{0}=E+i$,
$\operatorname{Im}\tilde{m}(z_{0})=-1-\operatorname{Im}m(z_{0})<-1.$
@@
For $k\geq 1$, we show that conditioning on $\Omega_{k-1}\cap\Xi_{k-1}$, the event $\Omega_{k}\cap\Xi_{k}$ holds with high probability. Define a threshold index $\widetilde{\mathcal{K}}:=\min\{k\in\mathbb{N}:k\leq\mathcal{K},|m(z_{k})-\tilde{m}(z_{k})|\leq 5\zeta_{k}t\}.$ For $k<\widetilde{\mathcal{K}}$, $|m(z_{k})-\tilde{m}(z_{k})|>5\zeta_{k}t$, and we must be sure that $s(z_{k})$ is close to $m(z_{k})$ rather than $\tilde{m}(z_{k})$. In contrast, for $k\geq\widetilde{\mathcal{K}}$, the distinction between $m(z_{k})$ and $\tilde{m}(z_{k})$ becomes insignificant, leading to a simpler argument. The majority of the proof follows steps analogous to those used for the base case $k=0$. We provide a sketch of the key arguments below.
@@ D.1 Proof of Corollary 1
$\mathbb{E}|X_{i}^{2}-\mathbb{E}X_{i}^{2}|^{l}\leq 2^{l}\mathbb{E}|X_{i}|^{2l}\leq 2^{l}p_{i}K^{2l}(2l)^{2l}.$ (D.1)